X+9=-2 WHERE DO I START WITH THIS PROBLEM? I NEED TO KNOW ALL THE SSTEPS I USE TO SOLVE FOR X. I KNOW THE ANSWER BUT DONT KNOW HOW TO GET IT USING A FORMULA.
Answer by longjonsilver(2297)

there is no formula. You just re-arrange the equation until you have x = a number and that is your answer.

x + 9 = -2 --> subtract 9 from both sides
x + 9 - 9 = -2 - 9
x = -11

this is the answer

jon.